# Privacy-Preserving in Connected and Autonomous Vehicles Through Vision to Text Transformation

### v2

@@ Abstract
Connected and Autonomous Vehicles (CAVs) rely on a range of devices that often process privacy-sensitive data. Among these, roadside units play a critical role particularly through the use of AI-equipped (AIE) cameras for applications such as violation detection. However, the privacy risks associated with captured imagery remain a major concern, as such data can be misused for identity theft, profiling, or unauthorized commercial purposes. While traditional techniques such as face blurring and obfuscation have been applied to mitigate privacy risks, individual privacy remains at risk, as individuals can still be tracked using other features such as their clothing. This paper introduces a novel privacy-preserving framework that leverages feedback-based reinforcement learning (RL) and vision-language models (VLMs) to protect sensitive visual information captured by AIE cameras. The main idea is to convert images into semantically equivalent textual descriptions, ensuring that scene-relevant information is retained while visual privacy is preserved. A hierarchical RL strategy is employed to iteratively refine the generated text, enhancing both semantic accuracy and privacy. Evaluation results demonstrate significant improvements in both privacy protection and textual quality, with the Unique Word Count increasing by approximately 77% and Detail Density by around 50% compared to existing approaches.
{IEEEkeywords}
Connected and Autonomous Vehicles, Feedback-based Learning, Privacy, Reinforcement Learning, Vision Language Model
@@ 1 INTRODUCTION
PRIVACY in Connected and Autonomous Vehicles (CAVs) has become a major concern in recent years, especially as a growing number of electronic devices are integrated into these vehicles. These systems transfer vehicle data to machine learning (ML)-based image processing pipelines for tasks such as image classification, object detection, face detection and recognition, image enhancement and restoration, and augmented reality. Among these technologies, Artificial Intelligence Equipped (AIE) cameras, which are increasingly deployed at junctions and traffic lights—are now integrated with CAV systems to monitor driver behavior (e.g., mobile phone use and seatbelt compliance) and to capture images for traffic monitoring, violation detection, and safety assessment, all aimed at enhancing public safety.
@@
Despite the financial and operational success of AIE camera systems, their deployment has raised serious privacy concerns. For example, in Queensland, Australia, an AIE camera initiative has generated approximately \$419.8 million in revenue since 2021 [1, 2], yet civil liberties advocates have objected to the use of these systems due to their potential to capture intimate images of drivers and passengers. As illustrated in Figure 1, the interior of vehicles—widely regarded as a private space—can be inadvertently exposed, raising legal and ethical concerns. Similar issues have been reported in U.S. states such as California and Maryland, where AIE cameras have been deployed despite ongoing legal challenges [3]. These deployments underscore broader concerns around visual data privacy in CAVs, which involve not only images but also GPS data, biometric identifiers, and audio streams [4].
Figure 1: Sample images captured in the Queensland project where the privacy is not preserved [1, 2].
@@
According to recent studies [5, 4], compromising the privacy of such data can result in consequences such as identity theft, unauthorized surveillance, user profiling, and commercial exploitation. Xie et al. [6] emphasize that the vehicle interior is an extension of personal space and must be protected accordingly. While techniques such as blurring and obfuscation are commonly used to mask sensitive content, they are often inaccurate and may fail to fully conceal private regions. Additionally, recent advances in AI-based reconstruction attacks pose serious threats, enabling adversaries to reverse these privacy measures and reveal the original content.
@@
Given the limitations of existing privacy-preserving techniques, such as inaccurate blurring and vulnerability to AI-based reconstruction attacks, there is a clear need for more robust solutions. In response, this study proposes a novel framework that transforms visual data into descriptive text using advanced VLMs, thereby avoiding the direct exposure of image content.
@@
The core idea of the proposed model is to leverage VLM techniques to generate descriptive text from collected images, capturing all necessary information without revealing the actual visual content. To enhance the quality and relevance of the generated descriptions, we introduce a feedback-based reinforcement learning (RL) framework that operates in parallel with a Retrieval-Augmented Generation (RAG) mechanism. RAG plays a key role in optimizing prompt selection by double-checking and refining the prompts used during RL training. To the best of our knowledge, this is the first attempt to combine RL and VLM for transforming visual data into text as a means of privacy preservation.
@@
RAG Retrieval Augmented Generation
MSE Mean Squared Error
@@
LLM Large Language Model
AIE AI-Equipped
MIPS Maximum Inner Product Search
@@ 2 RELATED WORK
There have been many privacy preservation studies where researchers have proposed different ways to manipulate parts of the target image. For example, in [10], the authors propose a face masking method to issue tickets. In another study [11] the authors focus on image obfuscation to protect privacy. Face obfuscation, masking, fogging, and replacements are the most common methods to avoid compromising privacy of images [12, 13]. However, existing methods are not capable to ensure privacy preservation. For example, Jian et al. in [14] explain how these methods are likely vulnerable to some AI methods such as adversarial attacks where attackers can recover identities. In a similar study [15], the author discusses the vulnerability of visual information hiding methods that are susceptible to adversarial attacks. In the Queensland case [1, 2], the goal was to detect driving violations, but parts of the driver’s body were still visible even though partial fogging is applied, and this raises additional privacy concerns. In a similar case, The State of Maryland uses AIE roadside cameras to detect traffic violations [3]. These cameras try to blur the driver’s face, but this does not fully guarantee visual privacy. In fact, a major problem with these methods is that they may incorrectly target parts of the image that involve privacy concerns [16, 6].
@@ 3 PROPOSED SOLUTION
This section presents the proposed hierarchical framework, which integrates VLM, RL, and RAG to address image privacy challenges in CAVs. Visual data is first captured by AIE cameras and transformed into textual descriptions using a VLM, which generates responses based on selected prompts aimed at minimizing exposure of sensitive content. A feedback-driven RL agent is employed to iteratively refine the prompt selection process. However, since the VLM’s prompt selection may not always yield optimal results, a RAG module is incorporated to monitor performance and enhance overall system effectiveness by retrieving and integrating relevant contextual information. The main components and operational flow of the proposed framework are detailed in the following four steps.
@@
At the heart of the model is a VLM model that converts images into textual descriptions. This component is particularly effective due to its dual-modality architecture, which enables the model to learn joint representations of visual and linguistic information, as illustrated in Figure  2. The VLM first performs object extraction with high accuracy, guided by the following loss function:
@@
Figure 2: Feedback-based RL-VLM model.
The VLM generates descriptions by matching images with textual prompts, allowing it to produce rich summaries of visual content. These textual prompts are not arbitrary but rather follow a hierarchical structured list of predefined prompts that guide the model in extracting key visual elements, such as vehicles, pedestrians, or existing conditions within the image. This ensures that the descriptions are relevant and actionable for further processing. By encoding complex images into text, this step compresses the data while maintaining the semantic information necessary for an accurate representation in the cloud environment.
Initially the model starts with token embedding $X=[x_{1},x_{2},...,x_{n}],x_{i}\in\mathbb{R}^{d}$ where $X$ is the input matrix of token vectors, $n$ is the length of the sequence, and $d$ is the embedding dimension. Positional encoding are added to represent token positions in a sequence because through the following equation since Transformers do not have a built-in notion of sequence order: $PE(pos,2i)=sin(\frac{pos}{10000^{\frac{2i}{d}}}),PE(pos,2i+1)=cos(\frac{pos}{10000^{\frac{2i}{d}}})$
@@
The final output is generated using the softmax function to convert scores into a probability distribution over the vocabulary. This is computed for each token to predict the next token in the sequence: $P(word_{i})=Softmax(z_{L}W^{V})$ where $W^{v}$ is the learned weight matrix that projects the final token embeddings into vocabulary size. To train the model, backpropagation is used to update the weights to minimize the loss. The key calculations here include cross-entropy Loss function and a gradient descent optimization through the following equations: $L=-\sum_{N}^{i=1}y_{i}log(\hat{y_{i}})$ and $\theta_{t+1}=\theta_{t}-\alpha\frac{m_{t}}{\sqrt{v_{t}}+\epsilon}$ where $m_{t+1}$ and $V_{t}$ are estimates of the first and second moments of the gradients, and $\alpha$ is the learning rate.
@@
In this step, initial textual descriptions generated by the VLM are refined using the proposed RL model. The RL model optimizes the text by mapping it to the most relevant and specific prompts from a refined list. This process enhances both relevance and specificity, resulting in concise yet context-aware descriptions that align with the requirements of the target digital environment.
@@
4:$i\leftarrow 1$
5:$iteration\leftarrow 3$
6:$feedback\leftarrow\textbf{False}$
7:$PU\leftarrow\textbf{None}$
@@
10:     if $feedback$ then
11:         UpdatePolicy($PU$)
@@
17:     EvaluationAndSelection($Text$, $Prompt\ List$)
@@
[…]
22:     if $PU$ then:
@@
Despite the successful extraction of the relative text from the images, the efficiency of the proposed model can be improved by adding an external function. The main idea is to propose a feedback-based RL framework to validate the process after extracting the refined text. To achieve this goal, RAG [19] employed as an external knowledge augmentation mechanism to evaluate the model performance based on the generated feedback. Integration RAG into the RL agent enables the proposed model to not only provide context-rich feedback but also improve the prompt management process, thereby improving the accuracy of the text refinement.
RAG offers advantages over PPO by dynamically retrieving the most relevant prompts, ensuring contextual accuracy without the need for retraining. In contrast, PPO relies solely on RL and may struggle with new or evolving data. Moreover, RAG is inherently more scalable and can efficiently operate over large knowledge bases.
To improve the accuracy of the generated text, the RAG model analyzes both the generated output and target prompts to identify the most relevant matches. This process occurs in parallel with the operation of the RL model. Once the RL agent selects the most relevant prompts, the RAG output is used to evaluate and validate the results. This stage is designed to correct potential errors or deficiencies in the RL model’s output, thereby supporting improved convergence and overall performance.
The feedback-based process creates a loop where the model iteratively improves its text generation by learning from these evaluations, enabling it to produce highly relevant and precise descriptions. The feedback mechanism operates based on the following equation:
@@
where $F(s,a)$ represents an external feedback function, and $\lambda$ is a weighting coefficient that determines the influence of the feedback on the learning process.
@@
The augmented RL objective combines the classic policy gradient approach—driven by the advantage term—with knowledge-driven external feedback. Specifically, $F(s_{t},a_{t})$ can represent a semantic similarity score, such as one computed by RAG between the generated text and the ground truth. This integration enables the agent to learn not only from environment-based rewards but also from external, domain-specific signals. The hyperparameter $\lambda$ governs the trade-off between traditional policy learning and feedback correction, allowing flexible adjustment based on the quality and availability of external knowledge. This formulation improves the model’s adaptability by integrating corrective feedback alongside environment-driven learning.
The RAG model’s evaluation process is based on Maximum Inner Product Search (MIPS), where the most relevant document $\hat{D}$ is retrieved by computing $\hat{D}=argmaxSim(q,D_{i})$, with $q$ denoting the query embedding generated by the encoder and $D_{i}$ representing the embedding of the $i-th$ document.

### v3

@@ Abstract
Intelligent Transportation Systems (ITS) rely on a variety of devices that frequently process privacy-sensitive data. Roadside units are important because they use AI-equipped cameras to detect traffic violations in Connected and Autonomous Vehicles (CAV). However, although the interior of a vehicle is generally considered a private space, the privacy risks associated with captured imagery remain a major concern, as such data can be misused for identity theft, profiling, or unauthorized commercial purposes. Methods like face blurring reduce privacy risks, however individuals’ privacy can still be compromised. This paper introduces a novel privacy-preserving framework that leverages feedback-based reinforcement learning (RL) and vision-language models (VLMs) to protect sensitive visual information captured by AIE cameras. The proposed idea transforms images into textual descriptions using an innovative method while the main scene details are preserved and protects privacy. A hierarchical RL strategy is employed to iteratively refine the generated text, enhancing both semantic accuracy and privacy. Unlike prior captioning-based methods, our model incorporates an iterative reinforcement-learning cycle with external knowledge feedback which progressively refines privacy-aware text. In addition to qualitative textual metric evaluations, the privacy-based metrics demonstrate significant improvements in privacy preservation where SSIM, PSNR, MSE, and SRRA values obtained using the proposed method on two different datasets outperform other methods.
keywords:
Feedback-based Learning, Privacy, Reinforcement Learning, Vision Language Model, Connected and Autonomous Vehicles
journal: Telematics and Informatics Reports
[…]
@@ 1 Introduction
Privacy in Intelligent Transportation Systems (ITS) has become a major concern in recent years, especially as a growing number of electronic devices are integrated into these vehicles. These systems transfer vehicle data to machine learning (ML)-based image processing pipelines for tasks such as image classification, object detection, face detection and recognition, image enhancement and restoration, and augmented reality. Among these technologies, Artificial Intelligence Equipped (AIE) cameras, which are increasingly deployed at junctions and traffic lights—are now integrated with CAV systems to monitor driver behavior (e.g., mobile phone use and seatbelt compliance) and to capture images for traffic monitoring, violation detection, and safety assessment, all aimed at enhancing public safety.
@@
Despite the financial and operational success of AIE camera systems, their deployment has raised serious privacy concerns. For example, in Queensland, Australia, an AIE camera initiative has generated approximately \$419.8 million in revenue since 2021 [10, 9], however potential privacy related issues arise from captured images. As illustrated in Figure 1, the interior of vehicles—widely regarded as a private space—can be inadvertently exposed, raising legal and ethical concerns. Similar issues have been reported in U.S. states such as California and Maryland, where AIE cameras have been deployed despite ongoing legal challenges [15]. These deployments highlight growing concerns about visual data privacy in connected and autonomous vehicles. The issue includes not only images but also GPS information, biometric data, and audio recordings [8].
Figure 1: Captured images are used to detect violations; however, as it is obvious, the privacy is compromised.
@@
According to recent studies [2, 8], compromising the privacy of such data can result in consequences such as identity theft, unauthorized surveillance, user profiling, and commercial exploitation. Xie et al. [18] emphasize that the vehicle interior is considered as personal space and must be protected. While techniques such as blurring and obfuscation are commonly used to mask sensitive content, they are usually inaccurate and may fail to fully conceal private regions. Additionally, recent advances in AI-based reconstruction attacks show serious threats, enabling adversaries to reverse these privacy measures and reveal the original content.
@@
Considering the limitations of existing privacy-preserving techniques, such as inaccurate blurring and vulnerability to AI-based reconstruction attacks, there is a need for more robust solutions. In response, this study proposes a novel framework that transforms visual data into descriptive text using advanced VLMs, thereby avoiding the direct exposure of image content.
Unlike traditional image captioning or single-pass text generation, our approach employs a hierarchical RL loop that iteratively refines semantic richness through external feedback from Retrieval Augmented Generation (RAG). Each iteration improves the contextual precision of descriptions beyond what static captioning or one-time VLM generation can achieve.
@@
The core idea of the proposed model is to leverage VLM techniques to generate descriptive text from collected images which captures all necessary information without revealing the actual visual content. To enhance the quality and relevance of the generated descriptions, we introduce a feedback-based Reinforcement Learning (RL) framework that operates in parallel with a RAG mechanism. RAG plays a key role in optimizing prompt selection by double-checking and refining the prompts used during RL training. To the best of our knowledge, this is the first attempt to combine RL and VLM for transforming visual data into text as a means of privacy preservation.
@@
ITS Intelligent Transportation Systems
LLM Large Language Model
MIPS Maximum Inner Product Search
MSE Mean Squared Error
@@
RL Reinforcement Learning
@@
VLM(s) Vision Language Model(s)
2 Background
Feedback-based RL is a specialized form of reinforcement learning where an agent learns through continuous feedback from the environment, but with an added emphasis on using richer, more informative feedback signals. Traditional RL typically relies on a simple reward signal, which might be sparse or delayed, meaning that the agent only receives feedback in the form of a reward or punishment based on the cumulative outcome of its actions. However, feedback-based RL introduces the concept of providing more nuanced or structured feedback, such as expert annotations, preferences, or evaluations from human users, to guide the agent toward better decision-making. This feedback could come from various sources and may not be limited to just numerical rewards; instead, it might include qualitative assessments, rankings, corrections, or other types of informative signals that help the agent understand what constitutes a ”good” or ”bad” action.
2.1 Mechanisms in Feedback-Based RL
In feedback-based RL, the agent still interacts with the environment by observing states, taking actions, and receiving feedback. However, the feedback may be more sophisticated and it often helps the agent to understand not just the outcome of its actions but also how to improve its decisions. For example, in a traditional RL setting, if an agent is learning to play a video game then it might only receive feedback when it wins or loses the game which can be sparse during the learning process. In contrast, feedback-based RL would involve providing the agent with more feedback such as pointers on specific decisions it makes, hints from an expert player, or even visual and verbal cues from the environment about how well it is progressing.
One of the key benefits of feedback-based RL is that it accelerates the learning process by reducing the reliance on sparse or delayed rewards. In environments where actions only produce feedback after a long sequence of decisions, traditional RL can struggle because the agent finds it difficult to attribute success or failure to specific actions. By incorporating more frequent and informative feedback, the agent is able to learn more efficiently. The feedback mechanism would be through the following ways:
  1. 1.
    Human Feedback: In some systems, humans provide feedback to guide the agent’s learning. This feedback can be explicit such as ranking actions or implicit feedbacks such as preference-based feedback where humans select actions they prefer.
  2. 2.
    Expert Annotations: In certain domains like healthcare or robotics, expert annotations can be used to guide agents. Experts can provide corrective actions or evaluations which help the agent to better understand the situation in which it is dealing with.
  3. 3.
    Preference Learning: In some cases, feedback comes in the form of preferences where an agent is given feedback comparing two different sequences of actions or outcomes which helps it understand which sequence leads to better outcomes.
  4. 4.
    Corrective Feedback: In robotics or control systems, an expert might directly intervene to correct the agent’s actions. This immediate, direct feedback helps the agent to refine the learning more quickly than waiting for the final outcome of a task.
Feedback-based RL is particularly useful in domains where high-quality or real-time feedback is available but learning from sparse, long-term rewards would be inefficient. By leveraging external feedback sources, the agent can often reach optimal policies faster and with less exploration.
2.2 Types of Feedback in Feedback-Based RL
Feedback-based RL can utilize different types of feedback beyond just rewards and penalties. One commonly used feedback type is human-in-the-loop feedback in which a human provides guidance to the agent in real time. This type of feedback can be useful in domains where an expert can offer insights that are not easily encoded into traditional reward functions. For example, in healthcare, a doctor can provide feedback on the quality of a treatment process devised by an RL agent, adjusting the agent’s decisions based on nuanced clinical knowledge. This expert feedback is better than a simple reward and provides the agent with actionable information about how to improve its policy.
Another form of feedback in this context is preference-based feedback, where an agent learns from feedback which involves ranking or comparing actions or outcomes. Instead of receiving a scalar reward for each action, the agent is presented with pairs of outcomes and asked which one is better. This is a more nuanced form of learning because the agent can start to understand relative preferences rather than learning purely from binary success or failure signals. This method has been particularly successful in applications like recommendation systems in which user’s preferences are expressed in comparative rather than absolute terms.
In addition, feedback can also be interactive, where an agent learns from an interaction with users or experts. This interaction allows the agent to clarify its understanding of the task or the environment and to receive specific feedback that helps it learn faster. For example, in robotics, an agent might ask a human operator for clarification when it encounters an unfamiliar object or a difficult task. By integrating feedback in real-time, the agent can improve its learning process dynamically and adjust its actions accordingly.
2.3 Applications of Feedback-Based RL
Feedback-based RL has numerous applications across different fields, particularly in areas where obtaining detailed feedback is possible and beneficial. One prominent application is in robotics, where agents can learn complex behaviors by receiving corrective feedback from human operators or experts. For example, a robot might be trained to grasp objects more effectively through a combination of expert guidance and real-time feedback, helping the agent refine its motor control skills.
In gaming, feedback-based RL has been used to improve agents that learn from human play or expert demonstrations. For instance, in games like StarCraft II or Dota 2, RL agents have been shown to perform better when they are guided by human experts who provide feedback on strategy, timing, and decision-making. By learning from expert strategies and receiving feedback on intermediate actions, the agents can achieve higher performance more quickly than through traditional RL methods alone.
Another area in which feedback-based RL excels is in recommendation systems, such as those used by streaming services or e-commerce platforms. Here, feedback comes in the form of user preferences or ratings, and the agent learns to optimize its recommendations based on what users prefer. For instance, instead of just presenting a set of items and learning from whether the user clicks or not, feedback-based RL can use feedback which users rank options or express preferences between different items. This feedback helps the agent understand more about the user’s underlying preferences, leading to better and more personalized recommendations.
In healthcare, feedback-based RL is used in personalized treatment planning. Doctors can provide feedback on the effectiveness of certain treatment actions, guiding the agent in optimizing treatment plans that are tailored to individual patients. By incorporating expert clinical knowledge through feedback, RL agents can learn to make more informed decisions, particularly in complex and high-stakes environments.
2.4 RL-based Captioning vs. Feedback-Driven Iterative Description
Traditional RL-based captioning methods focus primarily on optimizing a language model to generate a single caption that maximizes a linguistic or perceptual reward such as BLEU, CIDEr, or semantic similarity—based on one round of interaction between the RL agent and the captioning model. In these approaches, reinforcement learning serves mainly as a post-processing optimization step, fine-tuning captions to align with ground-truth descriptions. However, these models lack awareness of privacy, contextual feedback, or iterative refinement; once the caption is produced, the process ends without feedback-based adjustment or external evaluation of privacy preservation.
In contrast, the proposed feedback-driven iterative description framework extends beyond static RL captioning by introducing a multi-stage refinement loop where each cycle learns from both internal rewards (semantic similarity via SBERT) and external knowledge-based feedback (via RAG). This dual feedback mechanism enables the model to improve description quality meanwhile privacy is preserved across iterations. Instead of producing a single caption, the model evolves its descriptions over multiple rounds—filtering redundant content, enhancing semantic density, and ensures privacy preservation dynamically. This approach transforms RL from a one-time caption optimizer into a continuously adaptive reasoning agent.
@@ 3 Related Work
There have been many privacy preservation studies where researchers have proposed different ways to manipulate parts of the target image. For example, some authors propose a face masking method to issue tickets. In a different way, authors in [12] focus on image obfuscation to protect privacy. Face obfuscation, masking, fogging, and replacements are the most common methods to avoid compromising privacy of images [7, 5]. However, existing methods are not capable to ensure privacy preservation. For example, Jian et al. in [6] explain how these methods are likely vulnerable to some AI methods such as adversarial attacks where attackers can recover identities. In a similar study [20], the author discusses the vulnerability of visual information hiding methods that are susceptible to adversarial attacks. In the Queensland case [10, 9], the goal was to detect driving violations, but parts of the driver’s body were still visible even though partial fogging is applied, and this raises additional privacy concerns. In a similar case, The State of Maryland uses AIE roadside cameras to detect traffic violations [15]. These cameras try to blur the driver’s face, but this does not fully guarantee visual privacy. In fact, a major problem with these methods is that they may incorrectly target parts of the image that involve privacy concerns [17, 18].
Privacy in the context of CAVs is a complex term with multiple implications. For example, the authors in [3] focus on counting passengers at bus station without showing their faces. In [4], the goal is to collect traffic data while protecting individuals’ privacy. Another study by Zhang et al. [21] looks inside the vehicle to count passengers, also considering privacy concerns. Therefore, it can be said that the definition of privacy can vary based on the situation. There have been many privacy preservation studies where researchers have proposed different ways to manipulate parts of the target image. For example, some authors propose a face masking method to issue tickets. In a different way, authors in [12] focus on image obfuscation to protect privacy. Face obfuscation, masking, fogging, and replacements are the most common methods to avoid compromising privacy of images [7, 5].
Popescu et al. [15] proposed an obfuscation algorithm that combines a Variational Autoencoder (VAE) with random non-bijective pixel intensity mapping to protect sensitive content in medical images while maintaining sufficient fidelity for deep learning–based analysis. Their approach balances the privacy–accuracy trade-off by preventing human and AI-based reconstruction while enabling model training on obfuscated data. This method demonstrated strong performance for privacy-preserving classification in medical imaging, highlighting the potential of hybrid generative–statistical obfuscation strategies for sensitive visual data.
Jiang et al. [6] introduced DartBlur, a de-artifact blurring framework that uses a U-Net architecture and adversarial optimization to anonymize facial data while suppressing detection artifacts. Unlike traditional Gaussian blurring, DartBlur preserves operational fidelity and post-hoc fidelity which ensures consistent model performance on blurred and clean data. The authors also introduced multiple training objectives to enhance review convenience, accessibility and artifact suppression in order to represent a significant advancement in learnable blur-based privacy preservation for vision models.
Huang et al. [5] presented a dual surrogate generative model framework for identity-preserving face swapping that introduces explicit supervision via proxy-paired data. Their Credible Supervision Completion via Surrogates system creates synthetic source–target pairs using two surrogate networks to maintain identity features while modifying non-identity attributes. This dual-surrogate strategy improves visual fidelity and identity consistency compared to prior GAN or latent-based models which establishes a new direction for privacy-aware face synthesis.
Ghaziamin et al. [3] developed a privacy-preserving passenger counting system using overhead fisheye cameras and NVIDIA Jetson edge devices which enables real-time inference without cloud transmission. By performing all computations locally, the approach ensures user anonymity and reduces energy consumption. The system employs YOLO-based object detection optimized for fisheye distortion and demonstrates the practicality of responsible AI and edge computing for privacy-sensitive applications.
In Table 2 the main models are summarized and compared based on ten different features.
Table 2: Capability Comparison Among Privacy-Preserving Models
Capability [12] [6] [5] [3] Proposed model
Operates in Image Space $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
Operates in Semantic/Textual Space $\times$ $\times$ $\approx$ $\times$ $\checkmark$
Face/Identity Protection $\approx$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Generalization to Non-Facial Data $\checkmark$ $\times$ $\times$ $\approx$ $\checkmark$
Adversarial Robustness $\approx$ $\checkmark$ $\approx$ $\approx$ $\checkmark$
Model Interpretability $\times$ $\times$ $\approx$ $\approx$ $\checkmark$
Privacy–Utility Balance $\checkmark$ $\checkmark$ $\approx$ $\approx$ $\checkmark$
Human Review Convenience $\approx$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Scalability to Diverse Domains $\approx$ $\times$ $\approx$ $\approx$ $\checkmark$
Formal Privacy Guarantee (e.g., DP or abstraction) $\times$ $\times$ $\times$ $\times$ $\checkmark$
@@ 4 Proposed Solution
Figure 2: (a) General models follow a straightforward process which captures an image performing the image processing task and then produces the output. (b) The proposed architecture begins with image capture and generates textual descriptions through an iterative cycle, using new prompts in each iteration.
This section presents the proposed hierarchical framework which integrates VLM, RL, and RAG to address image privacy challenges in ITS. Existing devices capture and process images through a linear pipeline. In contrast, we propose an iterative process which combines VLM and RL, supported by a feedback mechanism that generates new textual descriptions at each iteration. More specifically, visual data is first captured by AIE cameras and transformed into textual descriptions using a VLM, which generates responses based on selected prompts aimed at minimizing exposure of sensitive content. A feedback-driven RL agent is employed to iteratively refine the prompt selection process. However, since the VLM’s prompt selection may not always generate optimal results, a RAG module is incorporated to monitor performance and enhance overall system effectiveness by retrieving and integrating relevant contextual information. The main components of the proposed framework are elaborated in the following four steps.
Figure 3: Feedback-based RL-VLM model running on AIE devices to capture images and transform into textual description through in an iterative process to generate maximum descriptions
@@
At the heart of the model is a VLM model that converts images into textual descriptions. This component is particularly effective due to its dual-modality architecture, which enables the model to learn joint representations of visual and linguistic information, as illustrated in Figure  3.
Unlike standard captioning models, the VLM is fine-tuned using privacy-sensitive supervision. Images are paired with privacy-annotated captions (e.g., ”driver visible”, ”face obscured”) and penalized when text exposes private details. A privacy classifier guides gradient updates so the VLM learns to avoid personally identifiable information (PII) descriptors during generation. This adaptation aligns the latent visual–text representations toward privacy-aware semantics.
The VLM first performs object extraction with high accuracy, guided by the following loss function:
@@
The VLM module generates descriptions by processing images based on prompts which allows it to produce rich summaries of visual content. These textual prompts are not arbitrary but rather follow a hierarchical structured list of predefined prompts that guide the model in extracting key visual elements, such as vehicles, pedestrians, or existing conditions within the image. This ensures that the descriptions are relevant and actionable for further processing. By encoding complex images into text, this step compresses the data while maintaining the semantic information necessary for an accurate representation in the cloud environment.
Initially the model starts with token embedding $X=[x_{1},x_{2},...,x_{n}],x_{i}\in\mathbb{R}^{d}$ where $X$ is the input matrix of token vectors, $n$ is the length of the sequence, and $d$ is the embedding dimension. Positional encoding are added to represent token positions in a sequence because through the following equation since Transformers do not have a built-in notion of sequence order: $PE(pos,2i)=sin(\frac{pos}{10000^{\frac{2i}{d}}}),PE(pos,2i+1)=cos(\frac{pos}{10000^{\frac{2i}{d}}})$.
@@
[…]
Figure 4: Hierarchical prompt selection where each layer prompts are selected at each iteration. As stated in Figure 3, Step 3, the generated text is utilized to select the most relevant prompts from the prompt list during each iteration, where the list is updated in every cycle.
@@
In this step, initial textual descriptions generated by the VLM are refined using the proposed RL model. While RL has been used in caption optimization, our contribution lies in integrating dual reward channels and hierarchical prompt refinement that iteratively enhance text relevance and privacy protection, an unexplored combination in current VLM literature. The RL model optimizes the text by mapping it to the most relevant and specific prompts from a refined list. This process enhances both relevance and specificity which results in concise and context-aware descriptions that align with the requirements of the target digital environment. As shown in Figure 4, the model initially selects the most relevant prompts from list A, then performs the same process with list B, and finally with list C. This iterative selection continues as the generated text guides the subsequent levels of the hierarchy.
@@
[…]
@@
10:  if feedbackfeedback then
11:   UpdatePolicy(PUPU)
@@
17:  EvaluationAndSelection(TextText, PromptListPrompt\ List)
@@
21:  $PU\leftarrow\text{PolicyUpdates(s)}$
22:  if PUPU then:
@@
Despite the successful extraction of the relative text from the images, the efficiency of the proposed model can be improved by adding an external function. The main idea is to propose a feedback-based RL framework to validate the process after extracting the refined text. To achieve this goal, RAG [1] is employed as an external knowledge augmentation mechanism to evaluate the model performance based on the generated feedback. Integrating RAG into the RL agent enables the proposed model to not only provide effective feedback but also it improves the prompt management process, thereby improving the accuracy of the text refinement.
RAG offers advantages over PPO by dynamically retrieving the most relevant prompts which ensures contextual accuracy without the need for retraining. In contrast, PPO relies solely on RL and may struggle with new or evolving data. Moreover, RAG is more scalable and can efficiently operate over large knowledge bases.
To improve the accuracy of the generated text, the RAG model analyzes both the generated output and target prompts to identify the most relevant matches. This process occurs in parallel with the operation of the RL model. Once the RL agent selects the most relevant prompts, the RAG output is used to evaluate and validate the results. This stage is designed to correct potential errors or deficiencies in the RL model’s output and therefore supports improved convergence and overall performance.
The feedback-based process creates a loop in which the model iteratively improves its description generation by learning from these evaluations which enables it to produce highly relevant and precise descriptions. The feedback mechanism operates based on the following equation:
@@
where $F(s,a)$ represents an external feedback function and $\lambda$ is a weighting coefficient that determines the influence of the feedback on the learning process.
@@
The augmented RL objective combines the classic policy gradient approach with external feedback. Specifically, $F(s_{t},a_{t})$ can represent a semantic similarity score such as one computed by RAG between the generated text and the ground truth. This integration enables the agent to learn not only from environment-based rewards but also from external, domain-specific signals. The hyperparameter $\lambda$ governs the trade-off between traditional policy learning and feedback correction, allowing flexible adjustment based on the quality and availability of external knowledge. This formulation improves the model’s adaptability by integrating corrective feedback alongside environment-driven learning.
The RAG model’s evaluation process is based on Maximum Inner Product Search (MIPS), where the most relevant document $\hat{D}$ is retrieved by computing $\hat{D}=argmaxSim(q,D_{i})$, with $q$ denoting the query embedding generated by the encoder and $D_{i}$ representing the embedding of the ithi-th document.